3b) In a locality of 18,000 families, a random sample of 840 families was taken. Of these

840 families, 206 families were found to have a monthly income of Rs. 500 or less.

Give the confidence interval for the families having income Rs. 500 or less.


1
Expert's answer
2021-05-07T10:14:50-0400

Percentage of families that have income as 500 or less are 20% and 28%.


Let the proportion of families that have income as 500 or less be = p


Thus,

"p = \dfrac{x}{n}



 = \dfrac{206}{840}



 = 0.24"


"q = 1-p\n\n\n\n = 1 - 0.24 = 0.75"


Therefore,


Expected number of families that have income as 500 or less = 18000 × p

"= 18000 \u00d7 0.24\n\n\n\n= 4410"


"\text{ S.e of p} =\dfrac{p(1-p)}{n}



=\dfrac{ 0.24 \times 0.75 }{ 840}



= 0.014"


Thus, the most probable limit for p will be - "0.24 + 3 ( 0.014)"


"- 0.200 \text{ and } 0.289"


Therefore, the percentage of families that have income as 500 or less are 20% and 28%.